OpenStudy (anonymous):

Write a rule for the linear function in the table.x -1,0,3,6 f(x) -1,2,5,8

OpenStudy (campbell_st):

what are the values of f(x) increasing by...?

OpenStudy (anonymous):

3 I think

OpenStudy (campbell_st):

ok... so thats the correct... but based on the ordered pairs (-1, -1) (0, 2) (3 , 5) and (6, 8) I think your 1st value may be incorrect...

OpenStudy (campbell_st):

should it be -3..?

OpenStudy (campbell_st):

oops x = -3, f(x) = -1

OpenStudy (anonymous):

It is the table says x-3,0,3,6 f(x) -1,2,5,8

OpenStudy (campbell_st):

thats makes more sense... your question had a typo... the general form of a straight line is y = mx + b m = slope and b = y-intercept... do you know this..?

OpenStudy (anonymous):

Yes sorry about that

OpenStudy (anonymous):

No just learning it

OpenStudy (campbell_st):

ok... so the y intercept occurs when x = 0.... so from the information what do you think the intercept is..?

OpenStudy (anonymous):

3

OpenStudy (campbell_st):

nope... look at the table of values, when x = 0 what is f(x) =

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

great thats call the y-intercept... and its where the line cuts the vertical axis so your equation is now y = mx + 2 does that make sense...?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks for the help

OpenStudy (campbell_st):

ok... so just a little more work... one of your points is x=3, f(x) = 5 so you need to substitute them into f(x) = mx + 2 to find the value of m.. is that ok..?

OpenStudy (anonymous):

Yes I understand it now thanks so much

OpenStudy (campbell_st):

ok... so what value do you have for m..?

OpenStudy (anonymous):

0

OpenStudy (campbell_st):

not quite if I substitute x = 3 and f(x) = 5 you get 5 = 3m + 2 whats the value of m..?

OpenStudy (anonymous):

5

OpenStudy (campbell_st):

not quite... its 3 times something plus 2 is equal to 5... whats the something..?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

I'm sorry but I'm dumb in some parts of math

OpenStudy (campbell_st):

thats correct... so the equation you have is f(x) = x + 2 since m = 1 and you can check by substituting other values

OpenStudy (anonymous):

Ok